# The role of porosity in the transition to inertial regime in porous media flows

### v1

@@
Abstract In this work we investigate the fundamental physical mechanism of the transition from Darcy to inertial (Darcy-Forchheimer) regime in the flow through porous media from the perspective of the vortex formation and growth. We focus on investigating the influence of porosity on the tortuosity–Reynolds number relation for periodic stochastic and simple cubic systems by numerically solving the Navier-Stokes equations in the pore-scale. We observe that the tortuosity defined by integrals over the fluid volume exhibits similar behavior in both types of systems. We also show that the tortuosity based on averaging of the streamlines length diverges from the volume-integrated one when the inertia onset takes place. We show that the discrepancy between those two tortuosities at increasing Reynolds number carries information about the geometrical confinement within the porous sample and is largely governed by the growth of the volumes occupied by vortices and their kinetic energy. Our results shed new light on the fundamental mechanisms of the transition between Darcy and the inertial regime and bring a better understanding of the role of the porous media geometry in this transition.
@@ 1 Introduction
Many real-life flows through porous media occur at velocities at which the effects caused by fluid inertia cannot be omitted. Such scenarios are of much importance for gas extraction [1, 2], mixing and reactive processes [3], or flow through network-like porous media [4], e.g. the circulatory system [5, 6]. It is thus crucial from scientific and application point of view to understand well the physical principles of the transition of flow through porous media from the creeping flow (Darcy) regime to the inertial regime and the behavior of the flow within the latter. Of special interest are the relations between the pore-scale hydrodynamics and the effective (volume-averaged) hydrodynamic coefficients of a porous medium, which serve as a link connecting the fundamental fluid dynamical processes to the more practical quantities characterizing those systems.
@@
stating the proportionality between the forcing term (mean pressure gradient $-\langle\nabla P\ \rangle$ and/or body force $\rho\boldsymbol{g}$) and the superficial flow velocity $\boldsymbol{U}$. The constant $k_{0}$ is the (intrinsic) permeability, characteristic for an isotropic porous system as $-\langle\nabla P\rangle$ and $\rho\boldsymbol{g}$ vanish, and $\nu$ is the fluid kinematic viscosity. In the inertial regime, the contribution of the convective derivative in the pore-scale momentum transport becomes non-negligible. Thus, the full Navier-Stokes equations need to be considered, yielding the Darcy-Forchheimer law [10]
@@
where $|\boldsymbol{U}|$ is the magnitude of the superficial velocity, $\beta$ is the free coefficient and the porous medium is again assumed to be isotropic. The transition from the Darcy to the inertial regime is often associated with the significant reorganization of the velocity and pressure fields within the pore spaces. Among many aspects of such reorganization, it was observed that the fluid kinetic energy becomes more evenly distributed in the pore space [11], which has to do with significant mass flow through an increasing number of channels and confinement of more kinetic energy within the vortices [12, 13, 14]. Our recent studies on high- [12] and low-porosity [14] porous media suggest that the porosity is one of the factors determining the dynamics of the transition from Darcy to inertial regime. Nevertheless, the exact mechanism of how the pores geometry impacts the onset of inertia is still unknown.
@@
For the quantification of the inertial effects in porous media, a variety of measures are used. Fitting Eqs. (1) and (2) to the experimental data allows to asses the permeability and the inertial correction coefficient and to distinguish between various inertial subregimes [13, 15, 11]. The measurements of the hydrodynamic drag force exerted on the solid fraction of the porous medium can also be used as the inertia onset indicator [16, 17]. Another quantity, commonly used in the porous media studies is the tortuosity $T$, in certain cases related to the permeability through the Carman-Kozeny relation [18, 19] as $k_{0}\propto T^{-1}$. Although originally introduced as an effective parameter describing solely the porous medium geometry, the notion of tortuosity was later extended to encompass factors other than geometry [20], and even to non-hydrodynamic phenomena [21]. Due to this, in its nearly 100-year’s history, several definitions of the tortuosity has been proposed and the discussion on their range of applicability and relation to each other is still ongoing [22, 23, 24]. Majority of the hydrodynamic (hydraulic) tortuosity definitions define it as an average of the lengths of the streamlines obtained from the pore-scale velocity field. However, it was proven that one can calculate the tortuosity directly from the pore-scale velocity field, without the need to generate streamlines beforehand [25]. This observation proved to be useful in low-porosity systems where even at low superficial velocities recirculation zones are abundant and the automatic seeding of the streamlines poses a big challenge. However, this equivalence between the two approaches to calculate the tortuosity no longer holds in the inertial regime, where the impact of the recirculation zones on the pore-scale flow is much larger than in the Darcy regime [25]. This introduces even more ambiguity into the description and explanation of the inertia onset in porous media.
@@
The aim of this work is to study how the interaction between the pores geometry and the pore-scale velocity drive the transition of porous media flow from the Darcy to the inertial regime. We investigate this transition quantitatively via two definitions of the tortuosity – one formulated using volume integrals of the velocity field, and the other, using the velocity streamlines. We explain why their dependence on the superficial flow velocity changes with the porosity of the porous sample. We show that those discrepancies, although seemingly contradictory to each other, provide information about the interaction between the geometry of the pores and the hydrodynamics in the pore-scale. In particular, we highlight that the main factor responsible for the behavior of the tortuosity during the Darcy-inertial transition is the dynamics of the growth of the recirculation zones in terms of their size and kinetic energy.
@@ 2.1.1 Governing equations and boundary conditions
We solve the steady-state incompressible Navier-Stokes equations in the volume of fluid $\Omega$ fully saturating the space between the solid phase of porous samples of side length $L$, with Dirichlet and Neumann boundary conditions on the solid walls $\partial\Omega$ for the velocity $\boldsymbol{V}$ and pressure $P$, respectively, and periodic boundary conditions along all three Cartesian coordinates
$\begin{array}[]{rll}\nabla\cdot\boldsymbol{V}=&\!\!\!\!0&\text{in }\Omega,\\ \boldsymbol{V}\>(\nabla\cdot\boldsymbol{V})=&\!\!\!\!-\nabla\dfrac{P}{\rho}+\boldsymbol{g}+\nu\nabla^{2}\boldsymbol{V}&\text{in }\Omega,\\ \boldsymbol{V}=&\!\!\!\!\boldsymbol{0}&\text{on }\partial\Omega,\\ \boldsymbol{n}\cdot\nabla P=&\!\!\!\!0&\text{on }\partial\Omega,\\ \boldsymbol{V}(\boldsymbol{x})=&\!\!\!\!\boldsymbol{V}(\boldsymbol{x}+\{0,\pm L\}^{3}),\\ P(\boldsymbol{x})=&\!\!\!\!P(\boldsymbol{x}+\{0,\pm L\}^{3}),\end{array}$ (3)
where the acceleration is aligned with the $x$-axis, $\boldsymbol{g}=[g,0,0]$ and $\boldsymbol{n}$ is the local unit normal vector on the solid boundary. We adopt the boldcase notation for the vectors and refer to the Cartesian coordinates with indices $0,1,2$, e.g. $\boldsymbol{V}\equiv[V_{0},V_{1},V_{2}]$. We denote the magnitude of a vector with a non-boldcase font, e.g. $|\boldsymbol{V}|\equiv V$.
It is useful to define at this point the notion of the percolating volume $\Omega_{p}\subset\Omega$. It is the volume consisting of all the streamlines that connect either two of the periodic boundaries of the fluid domain and are continuous within the porous sample (further called the percolating streamlines). The volume consisting of all fluid points which are not in $\Omega_{p}$ will be called the recirculation (or vortex) volume $\Omega_{v}$, i.e. $\Omega_{v}=\Omega\>\backslash\>\Omega_{p}$. Such definition of $\Omega_{v}$ is consistent with the definition of a steady-state vortex as the maximum set of nested closed streamlines used in previous works [26, 27]. With such division of $\Omega$, the boundary between the percolating and the recirculation volume, $\partial\Omega_{pv}$ can be clearly defined. A schematic representation of the percolating and recirculating volumes and the boundary between them is shown in Fig. 1.
Figure 1: A schematic representation of the percolating volume ($\Omega_{p}$, red), recirculation volume ($\Omega_{v}$, green) and the boundary between the two ($\partial\Omega_{pv}$, yellow) in a flow through a periodic porous sample (obstacles depicted gray). The curves with arrows along their length represent the percolating and the recirculating streamline. The dashed edges of the large square represent the periodic boundary.
@@ 2.1.2 Stochastic porous medium
The two stochastic porous samples considered in this work are confined within a cubic periodic cell of side length $L=1$. We generate the geometry of the solid phase by dividing the periodic cell into $16^{3}$ cubes and subsequently marking each of them as solid with probability equal to the desired porosity $\phi$ of the sample. We consider values $\phi=0.7$ and $\phi=0.9$. The visualizations of the solid phase of the two samples are shown in Fig. 2.
To solve the governing equations in the stochastic samples we use Lattice Boltzmann Method [28, 29] implemented for the GPU execution using Template Numerical Library [30, 31]. We use D3Q19 velocity set for $\phi=0.7$ sample and D3Q27 for $\phi=0.9$ sample. Two-relaxation time (TRT) collision kernel [32] with the ‘magic parameter’ $\Lambda=1/4$ and the symmetric relaxation time $\tau^{+}=0.505$ is used, along with the lattice speed of sound $c_{s}=1/\sqrt{3}$. The discretization in space consists of $N=304$ lattice sites along each direction for $\phi=0.7$ sample and $N=160$ lattice sites along each direction for $\phi=0.9$ sample. The no-slip walls are implemented via the halfway-bounceback method [28, 29]. In the simulations with the highest Reynolds number, in the final timestep, the maximal value of either velocity component for either samples did not exceed $3.73\cdot 10^{-2}$ and the density variation $|1-\rho|$ did not exceed $1.73\cdot 10^{-3}$, all in lattice units.
@@ 2.1.3 Simple cubic porous medium
The two simple cubic (SC) porous samples considered in this work are confined within a cubic periodic cell of side length $L=1$. The spherical obstacle is placed in the center of the periodic cell. We consider the obstacle radii $r=0.6526$ and $r=0.062$, resulting in porosities $\phi=0.1$ and $\phi=0.999$, respectively. For the convenience of visualization and meshing, the actual computational geometry for the lower-porosity sample considers the spherical obstacles placed in each of the periodic cell’s corners, i.e. the periodic cell is shifted by a vector $[L/2,L/2,L/2]$, The visualizations of the solid and fluid phases of the two samples are shown in Fig. 3.
We solve the governing equations with openFOAM [33]. We use the SIMPLE (Semi-Implicit Method for Pressure-Linked Equations) algorithm [34] for the pressure-velocity coupling, implemented in simpleFOAM solver. The convergence criterion based on residuals of the velocity and pressure is set to the values of $10^{-5}$ and $10^{-4}$, respectively. The space discretization is generated by first creating the regular cubic grid with the discretization parameter $h=1/150$ using blockMesh and then adjusting it to the actual fluid volume geometry with snappyHexMesh.
Figure 3: Visualization of the geometries of the solid (gray) and fluid (blue, opaque) phases of the two simple cubic porous samples considered in this work. The porosities are $\phi=0.1$ (left) and $\phi=0.999$ (right). The black-edged cube denotes the boundaries of the periodic cell. In the case of lowe-porosity sample only a few chosen obstacles residing in the periodic cell were chosen for clarity.
@@ 2.2 The used definitions of the tortuosity
To quantify the inertial effects in the flow, the present work considers two definitions of tortuosity. The first one, $T_{s}$, is defined as [23, 25, 22]
[…]
@@
where $\tilde{\lambda}(\boldsymbol{x})$ is the normalized tortuosity of a streamline passing through a point $\boldsymbol{x}\in A\cap\Omega$
@@
$\tilde{\lambda}(\boldsymbol{x})=\frac{\lambda(\boldsymbol{x})}{L_{\text{ref}}}$ (6)
with $L_{\text{ref}}$ denoting the reference length, unless otherwise stated we use $L_{\text{ref}}=L$. We will refer to this definition as the streamilne-based tortuosity. The integration in Eq. (5) is performed over the intersection of a plane $A$ perpendicular to the mean flow direction and the fluid volume $\Omega$. As noted by [25], the contribution of the recirculation zones to $T_{s}$ vanishes, thus Eq. (5) can be integrated over the intersection of the plane perpendicular to the mean flow direction and the percolating volume of the fluid
[…]
@@
although in numerical evaluations of $T_{s}$ it is more convenient to place $A$ such that it intersects only the percolating volume. In this, and all further integrations, one can equivalently assume that $\boldsymbol{V}=\boldsymbol{0}$ inside the solid phase and include the obstacles in the integrations. $T_{s}$ integrates the tortuosities of streamlines seeded from $A$ with weighting proportional to the flux in the mean flow direction associated with each streamline. Such weighting considers the streamlines of large flux, and thus carrying more fluid mass in a unit time, as more important for the value of the tortuosity, which increases the impact of the preferential channels of the flow [11]. It was proven in [25] that when there are no recirculation zones in the flow, $T_{s}$ is equivalent to the other tortuosity considered in the present work
[…]
@@
where the integration goes over the whole fluid volume $\Omega$. We will refer to this definition as the volume-integrated tortuosity. The notation $I_{\alpha}[\>\cdot\>]$ is a shorthand for the integral of some quantity over the fluid volume or its subset, i.e.
[…]
@@
Out of the two discussed tortuosities, $T_{s},V_{0}$ and $T_{\Omega}$, the latter turns out to be significantly easier to calculate numerically. This is because of its locality – it does not make use of streamlines, which, due to numerical errors, can easily cross the boundary $\partial\Omega_{pv}$ between the percolating and recirculation volume during the integration along the fluid velocity field characteristics. This problem becomes especially pronounced in low-porosity porous media and at high Reynolds numbers, where a considerable portion of the fluid is occupied by the recirculation zones.
2.3 Post processing of the results
Generation of streamlines
[…]

### v2

@@
Abstract In this work, we investigate the fundamental physical mechanism of the transition from Darcy to inertial (Darcy-Forchheimer) regime in steady-state flows through porous media, with the focus on vortex formation. We investigate their influence on the tortuosity–Reynolds number relation during this transition for systems of various porosities. We do so by numerically solving the Navier-Stokes equations within the pore-scale of simple cubic systems and relating the observations made therein to stochastic systems of more complex geometry. We observe that the tortuosity defined by integrals over the whole fluid volume behaves similarly in both types of systems. At the same time, in simple cubic systems, the tortuosity based on averaging of the length of the streamlines diverges from the volume-integrated one when the inertia onset takes place. We show that the discrepancy between those two tortuosities at increasing Reynolds number carries information about the dynamics of the vortex growth in the system. We stipulate that those dynamics are directly governed by the porosity. Our results highlight the utility of various definitions of tortuosity as measures of inertia in porous media flows and explain the reasons for the differences between those definitions. This can lead to a more sensible choice of inertia indicators in more application-oriented problems.
@@ 1 Introduction
Many real-life flows through porous media occur at velocities at which the effects caused by fluid inertia cannot be omitted. Such scenarios are of much importance for gas extraction [1, 2], mixing and reactive processes [3, 4], or flow through network-like porous media [5], e.g. the circulatory system [6, 7]. It is thus crucial from the scientific and application point of view to well well understand the physical principles of the transition of flow through porous media from the creeping flow (Darcy) regime to the inertial regime and the behavior of the flow within the latter. Of special interest are the relations between the pore-scale hydrodynamics and the effective (volume-averaged) hydrodynamic coefficients of a porous medium, which serve as a link connecting the fundamental fluid dynamical processes to the more practical quantities characterizing those systems.
@@
stating the proportionality between the forcing term (mean pressure gradient $\langle\nabla P\ \rangle$ and/or body force $\rho\boldsymbol{g}$) and the superficial flow velocity $\boldsymbol{U}$. The constant $k_{0}$ is the (intrinsic) permeability characteristic of an isotropic porous system as $\langle\nabla P\rangle+\rho\boldsymbol{g}$ term vanishes, and $\nu$ is the fluid kinematic viscosity. In the inertial regime, the contribution of the convective derivative in the pore-scale momentum transport becomes non-negligible. Thus, the full Navier-Stokes equations need to be considered, yielding the Darcy-Forchheimer law [11, 12]
@@
where $|\boldsymbol{U}|$ is the magnitude of the superficial velocity, $\beta$ is the free coefficient and the porous medium is again assumed to be isotropic. The transition from the Darcy to the inertial regime is often associated with the significant reorganization of the velocity and pressure fields within the pore spaces. Among many aspects of such reorganization, it was observed that the fluid kinetic energy becomes more evenly distributed in the pore space [13], which has to do with significant mass flow through an increasing number of channels and confinement of more kinetic energy within the vortices [14, 15, 16]. Our recent studies on high- [14] and low-porosity [16] porous media suggest that the porosity is one of the factors determining the dynamics of the transition from Darcy to inertial regime. Nevertheless, the exact mechanism of how the geometry of the pores impacts the onset of inertia is still unknown.
@@
For the quantification of the inertial effects in porous media, a variety of measures are used. Fitting Eqs. (1) and (2) to the experimental data allows one to assess the permeability and the inertial correction coefficient and to distinguish between various inertial subregimes [15, 17, 13]. The measurements of the hydrodynamic drag force exerted on the solid fraction of the porous medium can also be used as the inertia onset indicator [18, 19]. Another quantity, commonly used in porous media studies, is the tortuosity $T$, in certain cases related to the permeability through the Carman-Kozeny relation [20, 21] as $k_{0}\propto T^{-1}$. Although originally introduced as an effective parameter describing solely the porous medium geometry, the notion of tortuosity was later extended to encompass factors other than geometry [22], and even to non-hydrodynamic phenomena [23]. Due to this, in its nearly 100-year history, several definitions of the tortuosity have been proposed, and the discussion on their range of applicability and relation to each other is still ongoing [24, 25, 26]. The majority of the hydrodynamic (hydraulic) tortuosity definitions define it as an average of the lengths of the streamlines obtained from the pore-scale velocity field. However, it was proven that one can calculate the tortuosity directly from the pore-scale velocity field, without the need to generate streamlines beforehand [27]. It becomes especially useful in low-porosity systems where even at low superficial velocities, recirculation zones are abundant and the automatic seeding of the streamlines poses a big challenge. However, this equivalence between the two approaches to calculate the tortuosity no longer holds in the inertial regime, where the impact of the recirculation zones on the pore-scale flow is much larger than in the Darcy regime [27]. This introduces even more ambiguity into the description and explanation of the onset of inertia in porous media.
@@
The aim of this work is to study the behavior of tortuosity during the transition from the Darcy to the inertial regime. We do so to explain how the interplay between the pores’ geometry and the pore-scale velocity drives this transition in porous media flows. To describe the transition quantitatively, we use two definitions of the tortuosity – one formulated using volume integrals of the velocity field, and the other, using the velocity streamlines. We explain why their dependence on Reynolds number changes with the porosity of the porous sample. We show that those differences provide information about the interaction between the geometry of the pores and the hydrodynamics in the pore-scale. In particular, we highlight that the main factor affecting the tortuosity during the Darcy-inertial transition is the growth of the recirculation zones in size and kinetic energy.
@@ 2.1.1 Governing equations and boundary conditions
We solve the steady-state mass conservation (Eq. (3a)) and Navier-Stokes (Eq. (3b)) equations for an incompressible fluid in the pore-scale of three-dimensional porous samples. We consider cubic samples of side length $L$. The fluid fully saturates the pore space $\Omega$. On the solid walls, $\partial\Omega_{\text{solid}}$, we impose Dirichlet boundary conditions for the velocity $\boldsymbol{V}$ (Eq. (3c)) and Neumann boundary conditions for the pressure $P$ (Eq. (3d)). We also impose periodic boundary conditions along all three Cartesian coordinates (Eqs. (3e) and (3f))
$\nabla\cdot\boldsymbol{V}$ $=0$ $\quad\text{in }\Omega,$ (3a)
$\boldsymbol{V}\>(\nabla\cdot\boldsymbol{V})$ $=-\nabla\dfrac{P}{\rho}+\boldsymbol{g}+\nu\nabla^{2}\boldsymbol{V}$ $\text{in }\Omega,$ (3b)
$\boldsymbol{V}$ $=\boldsymbol{0}$ $\text{on }\partial\Omega_{\text{solid}},$ (3c)
$\boldsymbol{n}\cdot\nabla P$ $=0$ $\text{on }\partial\Omega_{\text{solid}},$ (3d)
$\boldsymbol{V}(\boldsymbol{x})$ $=\boldsymbol{V}(\boldsymbol{x}+\{0,\pm L\}^{3})$ $\text{in }\Omega,$ (3e)
$P(\boldsymbol{x})$ $=P(\boldsymbol{x}+\{0,\pm L\}^{3})$ $\text{in }\Omega.$ (3f)
The acceleration is aligned with the $x$-axis, $\boldsymbol{g}=[g,0,0]$, and $\boldsymbol{n}$ is the local unit normal vector on the solid boundary. We adopt the boldcase notation for the vectors and refer to the Cartesian coordinates with indices $0,1,2$, e.g. $\boldsymbol{V}\equiv[V_{0},V_{1},V_{2}]$. We denote the magnitude of a vector with a non-boldcase font, e.g. $|\boldsymbol{V}|\equiv V$.
It is useful to define at this point the notion of the percolating volume $\Omega_{p}\subset\Omega$. It is the volume consisting of all the streamlines that connect any pair of the periodic boundaries of the fluid domain and are continuous within the porous sample (further called the percolating streamlines). The volume consisting of all fluid points which are not in $\Omega_{p}$ will be called the recirculation (or vortex) volume $\Omega_{v}$, i.e. $\Omega_{v}=\Omega\>\backslash\>\Omega_{p}$. Such a definition of $\Omega_{v}$ is consistent with the definition of a steady-state vortex as the maximum set of nested closed streamlines used in previous works [28, 29]. With such division of $\Omega$, the boundary between the percolating and the recirculation volume, $\partial\Omega_{pv}$ can be clearly defined. A schematic representation of the volumes and the boundaries considered in this work is shown in Fig. 1.
Figure 1: A schematic representation of the percolating volume ($\Omega_{p}$, red), recirculation volume ($\Omega_{v}$, green) and the boundary between the two ($\partial\Omega_{pv}$, yellow) in a flow through a periodic porous sample (obstacles depicted striped gray with their boundary $\partial\Omega_{\text{solid}}$ plotted with solid black line). The curves with arrows along their length represent the percolating and the recirculating streamlines. The dashed edges of the large square represent the periodic boundary
@@ 2.1.2 Stochastic porous medium
The two stochastic porous samples considered in this work are confined within a cubic periodic cell of side length $L=1$. We generate the geometry of the solid phase by dividing the periodic cell into $16^{3}$ cubes and subsequently marking each of them as fluid with probability equal to the desired porosity $\phi$ of the sample. We consider values $\phi=0.7$ and $\phi=0.9$ and obtain the actual porosities of $0.696$ and $0.897$, respectively. The visualizations of the solid phase of the two samples are shown in Fig. 2.
To solve the governing equations in the stochastic samples, we use Lattice Boltzmann Method (LBM) [30, 31] implemented for the GPU execution using Template Numerical Library [32, 33]. Due to the uniform discretization of space, LBM allows for simple treatment of complex boundaries and has been successfully used to model flows in porous media [34, 35, 36], We use D3Q19 velocity set for $\phi=0.7$ sample and D3Q27 set for $\phi=0.9$ sample, for better stability at higher velocities. Two-relaxation time (TRT) collision kernel [37] with $\Lambda=1/4$ and the symmetric relaxation time $\tau^{+}=0.505$ is used, along with the lattice speed of sound $c_{s}=1/\sqrt{3}$. The discretization in space consists of $N=304$ lattice sites along each direction for $\phi=0.7$ sample and $N=160$ lattice sites along each direction for $\phi=0.9$ sample. The no-slip walls are implemented via the halfway-bounceback method [30, 31]. In the simulations with the highest Reynolds number, in the final timestep, the maximal value of either velocity component for either sample did not exceed $3.73\cdot 10^{-2}$ and the density variation $|1-\rho|$ did not exceed $1.73\cdot 10^{-3}$, all in lattice units.
@@ 2.1.3 Simple cubic porous medium
The two simple cubic (SC) porous samples considered in this work are confined within a cubic periodic cell of side length $L=1$. We consider porosities $\phi=0.1$ and $\phi=0.999$ corresponding to the obstacle radii $r=0.6526$ and $r=0.062$, respectively. For the convenience of visualization and meshing, the computational geometry for the lower-porosity sample considers the spherical obstacles placed in each of the periodic cell’s corners, and for the higher-porosity sample, we place a single obstacle in the center of the domain (see Fig. 3).
For SC samples, we solve the governing equations with openFOAM [38]. We use the SIMPLE (Semi-Implicit Method for Pressure-Linked Equations) algorithm [39] for the pressure-velocity coupling, implemented in simpleFOAM solver. In contrast to LBM used with the stochastic samples, SIMPLE allows one to solve directly for the steady-state solution, shortening the calculation time. For $\phi=0.1$, the convergence criterion based on residuals of the velocity and pressure is set to the values of $10^{-5}$ and $10^{-4}$, respectively. All considered systems reach this convergence criterion, and the verification of the steady-state solution for the system of the highest flow rate is performed with pimpleFOAM. The space discretization is generated by first creating the regular cubic grid with the discretization parameter $h=1/195$ using blockMesh and then adjusting it to the actual fluid volume geometry with snappyHexMesh (see Fig. 4).
For $\phi=0.999$, the convergence criterion based on residuals of the velocity and pressure is set to the values of $10^{-5}$ and $10^{-4}$, respectively. Three systems of the highest flow rate did not reach this criterion; however, the transient simulations performed with pimpleFOAM allow them to reach the steady state, and those data are used in the subsequent analysis. The determination of the steady state in those cases is based on the relative difference between the current and the estimated final value of the $x_{0}$-component of velocity in the bulk and in the wake of the obstacle. In the final iteration, this difference does not exceed $10^{-3}$. The final values of the velocities were estimated from the fit of the exponential function to the values of those velocities in time. We note that the $\phi=0.999$ system becomes unsteady in the range of $Re_{k}\in\>(5\cdot 10^{3};10^{4}]$. The discretization for $\phi=0.999$ systems solved using simpleFOAM is created the same way as for the $\phi=0.1$ systems. For the three highest flow rates solved with pimpleFOAM, we first create a regular cubic grid with the discretization parameter $h_{\text{bulk}}=1/75$ using blockMesh and then adjust it to the actual fluid volume geometry with snappyHexMesh with local refinement ($h_{\text{min}}=1/150$) near the sphere and in its wake (see Fig. 4). For all transient simulations (for both porosities), we choose the timestep $\Delta t$ such that the maximal Courant number in the system is below unity.
Figure 3: Visualization of the geometries of the solid (gray) and fluid (blue, opaque) phases of the two simple cubic porous samples considered in this work. The porosities are $\phi=0.1$ (left) and $\phi=0.999$ (right). The black-edged cube denotes the boundaries of the periodic cell. In the case of the low-porosity sample, only a few chosen obstacles residing in the periodic cell are shown for clarity.
Figure 4: Visualization of the meshes of the two simple cubic porous samples considered in this work. The porosities are $\phi=0.1$ (two top subplots) and $\phi=0.999$ (two bottom subplots, for the uniform and refined discretization). The top-right subplot shows the magnified view of the top-left mesh. The black-edged cube denotes the boundaries of the periodic cell. Only $z<0.5$ half of each mesh is shown.
@@ 2.2 The used definitions of the tortuosity
To quantify the inertial effects in the flow, let us introduce three definitions of tortuosity, two of which will be used further in this work. They mainly make use of the streamlines, defined as curves $s$ tangent to the velocity field vectors at each of their points. Equivalently, for the steady state systems, a streamline can be defined as a path along which a massless particle is advected in the velocity field (see Fig. 5). Since the velocity field has only one value in each fluid point $\boldsymbol{x}\in\Omega$, the streamlines cannot cross each other, and thus, there exists a 1-to-1 mapping between the fluid points and the streamlines. In other words, a streamline can be identified uniquely by any of its points: $s\equiv s(\boldsymbol{x})$ for any $\boldsymbol{x}\in s(\boldsymbol{x})$. We will use this fact when introducing the definitions of tortuosities in the following part of this section.
Figure 5: Schematic representation of a streamline $s$ as a curve tangent to the fluid velocity field vectors (black arrows) in each of the streamline’s points (two chosen are shown here as black dots, $\boldsymbol{x}$ and $\boldsymbol{x^{\prime}}$). An infinitesimal increment along a streamline, $d\boldsymbol{s}$ (small blue arrow), is parallel to the local velocity vector, here normalized for convenience. Any of the streamline’s points identifies it uniquely ($s(\boldsymbol{x})\equiv s(\boldsymbol{x^{\prime}})$).
The first presented definition of the tortuosity, $T_{s}^{*}$, was originally defined as a weighted average of the tortuosities of all the streamlines in the system [25, 27, 24]
$T_{s}^{*}=\frac{\displaystyle\int\limits_{A\>\cap\>\Omega}d\boldsymbol{x}\>V_{0}(\boldsymbol{x})\tilde{\lambda}(\boldsymbol{x})}{\displaystyle\int\limits_{A\>\cap\>\Omega}d\boldsymbol{x}\>V_{0}(\boldsymbol{x})}.$ (5)
@@
where $\tilde{\lambda}(\boldsymbol{x})$ is the tortuosity of a streamline passing through a point $\boldsymbol{x}\in A\cap\Omega$
@@
$\tilde{\lambda}(\boldsymbol{x})=\dfrac{\lambda(\boldsymbol{x})}{L_{\text{ref}}},\quad\lambda(\boldsymbol{x})\equiv\!\!\!\!\!\!\!\!\int\limits_{\text{along }s(\boldsymbol{x})}\!\!\!\!ds$ (6)
with $\lambda(\boldsymbol{x})$ denoting the (curvilinear) length of the streamline, $ds$ the length of the infinitesimal increment along the streamline, and $L_{\text{ref}}$ denoting the reference length. Unless otherwise stated, we take the reference length to be equal to the length of the porous sample’s side, i.e., $L_{\text{ref}}=L$. The integration in Eq. (5) is performed over the intersection of a plane $A$ perpendicular to the mean flow direction and the fluid volume $\Omega$ (see Fig. 6, leftmost plot). Also, $A$ must be intersected by all the streamlines in the system. As it is ambiguous to tell the length of a looping streamline inside a recirculation zone, the integration in Eq. (5) is usually limited only to the intersection of the plane perpendicular to the mean flow direction and the percolating volume of the fluid (see Fig. 6, middle plot)
$T_{s}=\frac{\displaystyle\int\limits_{A\>\cap\>\Omega_{p}}d\boldsymbol{x}\>V_{0}(\boldsymbol{x})\tilde{\lambda}(\boldsymbol{x})}{\displaystyle\int\limits_{A\>\cap\>\Omega_{p}}d\boldsymbol{x}\>V_{0}(\boldsymbol{x})},$ (7)
@@
although in numerical evaluations of $T_{s}$ it is more convenient to place $A$ such that it intersects only the percolating volume. In the present work, we will use the definition from Eq. (7) and will refer to it as the streamilne-based tortuosity. In this, and all further integrations, one can equivalently assume that $\boldsymbol{V}=\boldsymbol{0}$ inside the solid phase and include the obstacles in the integrations. $T_{s}$ integrates the tortuosities of the percolating streamlines seeded from $A\cap\Omega_{p}$ with weighting proportional to the flux in the mean flow direction associated with each streamline. Such weighting considers the streamlines of large flux, and thus carrying more fluid mass in a unit time, as more important for the value of the tortuosity, which increases the impact of the preferential channels of the flow [13]. It was proven in [27] that when there are no recirculation zones in the flow, $T_{s}$ is equivalent to the third tortuosity considered in the present work
$T_{\Omega}=\frac{\displaystyle\int\limits_{\Omega}d\boldsymbol{x}\>V(\boldsymbol{x})}{\displaystyle\int\limits_{\Omega}d\boldsymbol{x}\>V_{0}(\boldsymbol{x})}=\frac{I[V]}{I[V_{0}]}.$ (8)
@@
where the integration goes over the whole fluid volume $\Omega$ (see Fig. 6, rightmost plot). We will refer to this definition as the volume-integrated tortuosity. The notation $I_{\alpha}[\>\cdot\>]$ is a shorthand for the integral of some quantity over the fluid volume or its subset, i.e.
$I[f]_{\alpha}=\int\limits_{\Omega_{\alpha}}d\boldsymbol{x}\>f(\boldsymbol{x}),\quad\alpha\in\{\cdot,p,v\}$ (9)
@@
Out of the two tortuosities $T_{s}$ and $T_{\Omega}$, the latter turns out to be significantly easier to calculate numerically. This is because of its locality – it does not make use of streamlines, which, due to numerical errors during their generation, can easily cross the boundary $\partial\Omega_{pv}$ between the percolating and recirculation volume during the integration along the fluid velocity field characteristics. This problem becomes especially pronounced in low-porosity porous media and at high Reynolds numbers, where a considerable portion of the fluid is occupied by the recirculation zones.
Figure 6: Schematic representation of the three definitions of tortuosities from Eqs. (5), (7), and (8). $T_{s}^{*}$ (left) and $T_{s}$ (middle) integrate over the surface elements $d\boldsymbol{x}$ of the intersections of an arbitrary surface AA and the corresponding fluid volumes. $T_{s}^{*}$ considers all the fluid volume $\Omega$, while $T_{s}$ only its percolating part, $\Omega_{p}$. $T_{\Omega}$ (right) integrates over the whole fluid volume $\Omega$. $T_{s}$ and $T_{\Omega}$ are used further in the present study.
2.3 Results analysis
2.3.1 Generation of streamlines
To generate the streamlines in SC samples, we numerically integrate the advection equation for the massless tracers advected in the fluid velocity field. We use the first-order explicit Euler method for the timestepping and meshless RBF-FD approximation [40] for the interpolation in space. Cubic radial functions augmented with second-order monomials are chosen as the interpolation basis, along with 25 closest neighbors in the approximation stencil. We use the C++ implementation of meshless approximation routines provided by Medusa library [41]. For the generation of streamlines, we use the normalized velocity field, $\boldsymbol{V}/V$, i.e., we set the magnitude of the velocity vectors to unity and preserve only their direction (see Fig. 7).
Figure 7: Schematic representation of the streamline generation procedure. The current increment of the streamline, $\Delta\boldsymbol{s}$ (denoted with an arrow), starts with the interpolation of the velocity to the current streamline end (blue filled circle). The interpolation uses a meshless scheme (and is thus naturally capable of operating on non-regular grids) and considers a number of the mesh points closest to the current streamline tip (filled squares, encircled in a dashed loop). The next streamline point (open circle) is calculated using the first-order Euler time integration.